# Evaluating Uncertainty-based Failure Detection for Closed-Loop LLM Planners

### v1

@@ Abstract
Recently, Large Language Models (LLMs) have witnessed remarkable performance as zero-shot task planners for robotic manipulation tasks. However, the open-loop nature of previous works makes LLM-based planning error-prone and fragile. On the other hand, failure detection approaches for closed-loop planning are often limited by task-specific heuristics or following an unrealistic assumption that the prediction is trustworthy all the time. As a general-purpose resoning machine, LLMs or Multimodal Large Language Models (MLLMs) are promising for detecting failures. However, the appropriateness of the aforementioned assumption diminishes due to the notorious hullucination problem. In this work, we attempt to mitigate these issues by introducing a framework for closed-loop LLM-based planning called KnowLoop, backed by an uncertainty-based MLLMs failure detector, which is agnostic to any used MLLMs or LLMs. Specifically, we evaluate three different ways for quantifying the uncertainty of MLLMs, namely token probability, entropy, and self-explained confidence as primary metrics based on three carefully designed representative prompting strategies. With a self-collected dataset including various manipulation tasks and an LLM-based robot system, our experiments demonstrate that token probability and entropy are more reflective compared to self-explained confidence. By setting an appropriate threshold to filter out uncertain predictions and seek human help actively, the accuracy of failure detection can be significantly enhanced. This improvement boosts the effectiveness of closed-loop planning and the overall success rate of tasks.
@@ III-A Uncertainty-based Failure Detection
To trigger an MLLM failure detector, a proper and informative prompt is necessary. For this purpose, we specifically design two representative categories of prompting strategies, namely a direct one and an indirect one. Direct prompting strategy entails directly the question of success or failure, while the indirect one does not encompace this kind of direct question of failure or success in the context. Furthermore, to tackle the improper assumption of the failure detector, we consequently estimate the uncertainty of each response to these prompts. Based on the uncertainty estimates and a predetermined threshold, the reliability of each response can be evaluated quantitatively in a generic way. Responses with excessively high uncertainty will be filtered out, and instead, human feedback will be requested to assess the current state, as shown in Algorithm 1. More noteworthy, the proposed way to detect failure cases is agnostic to LLM used in the planner. Therefore, it enjoys the merit of general applicability when compared with other task-specific or heuristic approaches.
III-A1 Prompting Strategy
@@
Direct Prompts via Subgoal State Comparison (SSC). Given only background information and environmental images, inferring the state description upon completing an action is challenging for LLMs. Moreover, comparing this with the environmental images to determine if the current state aligns with the subgoal poses additional challenges. Consequently, we propose to succinctly outline each subgoal or action, accompanied by an expected state description. The MLLM will then compare this description against the current environmental image to ascertain whether the present state matches the anticipated outcome. This indicates the successful achievement of the subgoal.
@@
Direct Prompts via Spatial Relationship Analysis (SRA). The ”Chain of Thought” (CoT) [32] reasoning in LLMs epitomizes a sophisticated strategy for tackling complex problem-solving tasks. This technique capitalizes on the model’s capacity to generate intermediate steps or rationales leading to a conclusive answer, thereby emulating a process akin to human reasoning. In the context of robot manipulation tasks, the most profound transformations are observed in the dynamics between the gripper and the spatial positioning of objects following each action. Rather than directly evaluating the outcome as success or failure based on state descriptions and images, the MLLM should incorporate an intermediate cognitive process. This analyzes the current spatial positions and conditions of objects within the environment, comparing these with the anticipated states and subsequently adjudicating the action’s success or failure.
@@
Indirect Prompts via Next Action Prediction(NAC). To distinguish this strategy from the preceding ones, success is no longer binary (a simple yes or no). Instead, upon completing each action, we present all high-level plans and the corresponding executable actions to the MLLM. Then, images are supplied to illustrate the current state of the environment. Based on the current state, the MLLM is asked to select the subsequent action from the presented options. This approach reformulates the detection problem into a multiple-choice scenario. Success is achieved when the MLLM’s selection aligns with the subsequent step of the predefined plan. Conversely, the selection of any alternative action is deemed an execution failure.
@@
III-A2 Uncertainty Quantification
The precision of the failure detector’s judgments is pivotal for the success of the entire closed-loop control system. Should a true negative scenario arise, it may cause the system to miss errors during execution, persisting with the initial plan and leading to task failure. In contrast, a false negative can interrupt the correct sequence of operations, leading to erroneous states that hinder the task’s completion. Hence, the greater the failure detector’s accuracy, the more effective the system. Nonetheless, even the most advanced MLLMs, like ChatGPT-4V, fail to achieve flawless accuracy and may produce hallucinations, generating responses arbitrarily. We aim to quantify the uncertainty in the model’s responses to better gauge its confidence in its answers. When the model is uncertain, expressing ”I don’t know” is preferable to providing an arbitrary response.
Token probability. Prior work [25] suggests that, by formulating multiple-choice questions and answering (MCQA) and having LLMs’ responses to them, it is feasible to deduce the occurrence probability of each option token $P(x_{i})$. This probability is interpreted as the likelihood of the corresponding option being correct. In the context of failure detection, the issue can be conceptualized as a binary question: whether the outcome of the most recent action was successful or if the current state aligns with the anticipated state. The response sought is either ’Yes’ or ’No’. In a similar vein, deriving the token probability from the MLLM output can act as an indicator of the probability of affirmative or negative responses.
Entropy. In information theory, entropy represents a measure of uncertainty or disorder. For binary classification problems, the entropy of a model’s predictions can be calculated using the formula:
$H(X)=-\sum_{i}P(x_{i})\log P(x_{i})$ (1)
where $P(x_{i})$ is the probability of class $i$ (in case of failure detection, ”Yes” or ”No”). Within the output layer of the MLLM, each token is assigned a score. When subjected to a softmax transformation, these scores yield the probability associated with each token. The probability of a token can thus be interpreted as the probability of a given classification. The entropy derived from these token probabilities may reflect the model’s uncertainty about its response to a certain extent. Higher entropy values suggest greater model uncertainty. This relationship between entropy and model uncertainty will be further explored and discussed in the subsequent sections.
@@ IV Experiment
We first show our experimental setup including the self-collected dataset and implementation details of the system and algorithm. Then we discuss the experiment results. We first evaluate the performance of three different uncertainty estimation methods along with different promoting strategies based on a self-collected dataset.
Secondly, based on the results obtained in the uncertainty evaluation part, we select the combination with the most encouraging performance for closed-loop LLM planners on real hardware. To note that the test set includes slightly different tasks [^1] Task_1:pick up the sponge and place it on the notebook, Task_2:pick up the smartphone and place it in the upper drawer, Task_3:open the upper drawer but with similar items. The LLM initiates high-level planning upon receiving language instructions, decomposing the tasks into several subtasks. In scenarios devoid of external interference, the robotic arm’s adherence to the high-level action plan results in a success rate ranging between 70-80% for these tasks. To illustrate the significance of closed-loop control more effectively, each experiment will introduce human interference—such as displacing objects from their intended locations or impeding operations. This interference leads to a success rate of 0% for these tasks under open-loop control conditions. In this experiment, we also compare ChatGPT-4V and LLaVA to show the general compatibility of our proposed framework.
@@ IV-B Evaluation Metrics
In this subsection, we explain the metrics used for evaluating different variants in the experiment.
Uncertainty-Acuracy Curve and Area Under Curve (AUC) (Calibration-AUC)
The uncertainty-accuracy curve [18] depicts accuracy as a function of the abstention uncertainty threshold. This threshold determines the point at which samples are excluded from consideration if their uncertainty surpasses $(1-threshold)$. Initially, no samples are omitted when the uncertainty threshold is set to zero, allowing the accuracy measurement to reflect the entire dataset’s performance. As the uncertainty threshold escalates, samples exhibiting uncertainty above $(1-threshold)$ are selectively discarded. Subsequently, the accuracy metric is recalculated to reflect the composition of the residual dataset. Should uncertainty serve as a reliable indicator of the model’s uncertainty, discarding samples marked by higher uncertainty should theoretically enhance the quality of the retained samples. Consequently, this selective exclusion is expected to improve the model’s overall accuracy.
Selective Generation Curve and AUC (Selective-AUC)
The selective generation curve [18] depits accuracy as a function of the abstention rate, denoted as $\alpha$. This approach involves ordering the samples by their uncertainty and then abstaining from the top $\alpha\%$ of samples based on their uncertainty values. No samples are abstained at $\alpha=0\%$, meaning the curve initiates at the conventional accuracy metric. Consequently, an increase in the curve is anticipated as $\alpha$ escalates.
Success Rate
For each real-world task, it is considered a success if the instruction requirements are eventually met under our closed-loop control. If the task remains incomplete after reaching the maximum number of attempts, it is deemed a failure. The success rate is calculated by dividing the number of successful experiments by the total number of experiments.
Detection Accuracy
Detection accuracy is the number of correct predictions divided by the total number of detections made by the MLLM failure detector. Instances where humans provide assistance are not included. This metric indicates the ability of MLLM to make accurate predictions.
Human Involve Rate
The metric is calculated by dividing the number of times humans provided assistance by the total number of detections performed during the experiment. A higher human involvement rate indicates a larger effort to achieve a high success rate, which further implies the limited performance of the MLLM failure detector.
Generation Rate
This metric only applies to Self-explained Confidence, as not every attempt results in a successful generation of confidence scores. It is calculated as the number of successful confidence score generations divided by the total number of tests.
@@ IV-D1 Uncertainty Estimation for MLLM Failure Detecor
We first evaluate the MLLM failure detector based on LLaVA, as we need token probability for uncertainty calculation which is not supported by ChatGPT-4V API at the moment. We document each sample’s prediction outcomes and token probability for each prompting strategy. Upon quantification, the utility of LLaVA’s entropy, as derived through the techniques mentioned above, in differentiating between high-quality and low-quality predictions remains ambiguous. Consequently, we employ the following criteria to assess the correlation between the uncertainty derived from various prompting and quantification methods and the accuracy of the predictions shown in Table II, Fig. 4 and Fig. 5.
@@
The increasing curves in Fig. 4 indicate that uncertainty quantification through token probability and entropy yields promising performance. A distinct trend emerges, showing increased accuracy as uncertainty diminishes. However, as observed in Fig. 5, the correlation between self-reported uncertainty and accuracy is weak, rendering the derived uncertainty measures impractical for use.
In Table II, the AUC values for the strategies of SSC and SRA are relatively similar and notably superior to that of the NAP strategy. This disparity is primarily attributed to the lower accuracy of the NAP strategy in the absence of filtering. The linear correlation between accuracy and entropy-based uncertainty, particularly evident in the SSC prompt, is the most pronounced among the measures examined.
For the promising performance of SSC with entropy, we adopt the curve produced by the SSC prompting strategy with entropy for the real hardware experiment.
@@ IV-D2 Uncertainty for Closed-Loop LLM Planners
In the real-world experiment, we chose ChatGPT-4V and LLaVA as baselines to compare with the proposed framework KnowLoop. In both experiments with these MLLMs, the detector fully trusts the MLLM’s prediction results without considering uncertainty. For KnowLoop, we filter the failure predictions based on uncertainty with a threshold of 0.6, above which the corresponding prediction will not be trusted, and the model will actively seek human help for failure detection. In Table III, as can be seen, when compared to the baselines that completely trust the MLLM’s predictions, our framework can, to some extent, improve the task success rate by around 50% compared to ChatGPT-4v and by around 40% compared to LLaVA. The promising results demonstrate that KnowLoop can enable the robot to actively ask for human help when the prediction is uncertain.
ing
Figure 5: Selective Generation Curve for three prompting strategies based on Self-explained Confidence.

### v2

@@ Abstract
Recently, Large Language Models (LLMs) have witnessed remarkable performance as zero-shot task planners for robotic manipulation tasks. However, the open-loop nature of previous works makes LLM-based planning error-prone and fragile. On the other hand, failure detection approaches for closed-loop planning are often limited by task-specific heuristics or following an unrealistic assumption that the prediction is trustworthy all the time. As a general-purpose resoning machine, LLMs or Multimodal Large Language Models (MLLMs) are promising for detecting failures. However, the appropriateness of the ”always true” assumption gets exacebated due to the notorious hullucination problem. In this work, we attempt to mitigate these issues by introducing a framework for closed-loop LLM-based planning called KnowLoop, backed by an uncertainty-based MLLMs failure detector, which is agnostic to any used MLLMs or LLMs. Specifically, we evaluate three different ways for quantifying the uncertainty of MLLMs, namely token probability, entropy, and self-explained confidence as primary metrics based on three carefully designed representative prompting strategies. With a self-collected dataset including various manipulation tasks and an LLM-based robot system, our experiments demonstrate that token probability and entropy are more reflective compared to self-explained confidence. By setting an appropriate threshold to filter out uncertain predictions and seek human help actively, the accuracy of failure detection can be significantly enhanced. This improvement boosts the effectiveness of closed-loop planning and the overall success rate of tasks.
@@ III-A Uncertainty-based Failure Detection
A proper and informative prompt is necessary to trigger an MLLM failure detector. For this purpose, we specifically design two representative categories of prompting strategies, namely a direct one and an indirect one. Direct prompting strategy directly entails the question of success or failure, while indirect one does not encompass this kind of direct question of failure or success in the context. Furthermore, to tackle the improper assumption of the failure detector, we consequently estimate the uncertainty of each response to these prompts. Based on the uncertainty estimates and a predetermined threshold, the reliability of each response can be evaluated quantitatively in a generic way. Responses with excessively high uncertainty will be filtered out, and instead, human feedback will be requested to assess the current state, as shown in Algorithm 1. More noteworthy, the proposed way to detect failure cases is agnostic to LLM used in the planner. Therefore, it enjoys the merit of general applicability compared to other task-specific or heuristic approaches.
III-A1 Prompting Strategies
@@
Direct Prompts via Subgoal State Comparison (SSC). Inferring the state after an action and comparing it to the environmental images is challenging for LLMs. To address this, we propose outlining each subgoal or action, along with its expected state description. The MLLM will then compare this with the current environmental image to determine if the present state aligns with the anticipated outcome, indicating the successful achievement of the subgoal.
Direct Prompts via Spatial Relationship Analysis (SRA). The ”Chain of Thought” (CoT) [32] reasoning in LLMs epitomizes a sophisticated strategy for tackling complex problem-solving tasks. The technique utilizes the model’s ability to generate intermediate steps or reasons, mimicking human reasoning. In robot manipulation tasks, notable changes occur in the interaction between the gripper and the spatial arrangement of objects after each action. Rather than solely assessing success or failure based on state descriptions and images, the MLLM should involve an intermediate cognitive process that compares current and expected states to determine the action’s outcome.
Indirect Prompts via Next Action Prediction(NAC). To differentiate this approach from previous ones, success is no longer a simple yes or no. Instead, after each action, we present all high-level plans and corresponding executable actions to the MLLM. We then provide images to illustrate the current environment. Based on this, the MLLM selects the next action from the options presented, turning the detection problem into a multiple-choice scenario. Success occurs when the MLLM’s selection aligns with the next step of the predefined plan, while any alternative action is considered a failure.
III-A2 Uncertainty Quantification
The failure detector’s accuracy is crucial for the closed-loop control system’s success. True negatives can cause the system to miss errors, leading to task failure, while false negatives can disrupt operations. Even advanced MLLMs like ChatGPT-4V may produce hallucinating answers, and it’s important to gauge their confidence. When uncertain, it’s better to say ”I don’t know” than provide a false response.
@@ III-A2 Uncertainty Quantification
Token probability. Prior work [25] suggests that, by formulating multiple-choice questions and answering (MCQA) and having LLMs’ responses to them, it is feasible to deduce the occurrence probability of each option token $P(x_{i})$. This probability is interpreted as the likelihood of the corresponding option being correct. Deriving the token probability from the MLLM output can act as an indicator of the probability of affirmative or negative responses.
Entropy. In information theory, entropy represents a measure of uncertainty or disorder. Within the output layer of the MLLM, each token is assigned a score. The probability of a token can thus be interpreted as the probability of a given classification. The corresponding entropy may reflect the model’s uncertainty about its response to a certain extent.
@@ IV Experiment
We will first present our experimental setup and implementation details, followed by a discussion of the experiment results. Then, we will evaluate three uncertainty estimation methods and select the best-performing combination for closed-loop LLM planners on real hardware based on the results. To note that the test set includes slightly different tasks [^1] Task_1:pick up the sponge and place it on the notebook, Task_2:pick up the smartphone and place it in the upper drawer, Task_3:open the upper drawer but with similar items. In interference-free scenarios, the robotic arm successfully completes tasks $70-80\%$ of the time. When human interference is introduced—such as displacing objects from their intended locations or impeding operations, the success rate drops to $0\%$ under open-loop control conditions.
@@ IV-B Evaluation Metrics
In this subsection, we explain the metrics used for evaluating different variants in the experiment: a) Uncertainty-Acuracy Curve and Area Under Curve (AUC) (Calibration-AUC) [18] depicts accuracy as a function of the abstention uncertainty threshold. This threshold determines the point at which samples are excluded from consideration if their uncertainty surpasses $(1-threshold)$. Initially, no samples are omitted when the uncertainty threshold is set to zero, allowing the accuracy measurement to reflect the entire dataset’s performance. As the uncertainty threshold escalates, samples exhibiting uncertainty above $(1-threshold)$ are selectively discarded. Subsequently, the accuracy metric is recalculated to reflect the composition of the residual dataset. Should uncertainty serve as a reliable indicator of the model’s uncertainty, discarding samples marked by higher uncertainty should theoretically enhance the quality of the retained samples. Consequently, this selective exclusion is expected to improve the model’s overall accuracy. b) Selective Generation Curve and AUC (Selective-AUC) [18] depits accuracy as a function of the abstention rate, denoted as $\alpha$. This approach involves ordering the samples by their uncertainty and then abstaining from the top $\alpha\%$ of samples based on their uncertainty values. No samples are abstained at $\alpha=0\%$, meaning the curve initiates at the conventional accuracy metric. Consequently, an increase in the curve is anticipated as $\alpha$ escalates. c) Success Rate is calculated by dividing the number of successful experiments by the total number of experiments. d) Detection Accuracy is the number of correct predictions divided by the total number of detections made by the MLLM failure detector. Instances where humans provide assistance are not included. This metric indicates the ability of MLLM to make accurate predictions. e) Human Involve Rate is calculated by dividing the number of times humans provided assistance by the total number of detections performed during the experiment. A higher human involvement rate indicates a larger effort to achieve a high success rate, which further implies the limited performance of the MLLM failure detector. f) Generation Rate only applies to Self-explained Confidence, as not every attempt results in a successful generation of confidence scores. It is calculated as the number of successful confidence score generations divided by the total number of tests.
@@ IV-D1 Uncertainty Estimation for MLLM Failure Detecor
We first evaluate the MLLM failure detector based on LLaVA, as we need token probability for uncertainty calculation which is not supported by ChatGPT-4V API at the moment. We employ the following criteria to assess the correlation between the uncertainty derived from various prompting and quantification methods and the accuracy of the predictions shown in Table II, Fig. 4.
The increasing curves in Fig. 4 indicate that uncertainty quantification through token probability and entropy yields promising performance. A distinct trend emerges, showing increased accuracy as uncertainty diminishes. However, the correlation between self-explained uncertainty and accuracy is weak, rendering this measure impractical for use. In Table II, the AUC values for the strategies of SSC and SRA are relatively similar and notably superior to that of the NAP strategy. This disparity is primarily attributed to the lower accuracy of the NAP strategy in the absence of filtering. The linear correlation between accuracy and entropy-based uncertainty is most noticeable in the SSC prompt. We have chosen to use the curve generated by the SSC prompting strategy with entropy for the real hardware experiment due to the promising performance of SSC with entropy.
@@ IV-D2 Uncertainty for Closed-Loop LLM Planners
In the real-world experiment, we chose ChatGPT-4V and LLaVA as baselines to compare with the proposed framework KnowLoop. In both experiments with these MLLMs, the detector fully trusts the MLLM’s prediction results without considering uncertainty. For KnowLoop, we filter the failure predictions based on uncertainty with a threshold of 0.6, above which the corresponding prediction will not be trusted, and the model will actively seek human help for failure detection. In Table III, when compared to the baselines that completely trust the MLLM’s predictions, our framework can, to some extent, improve the task success rate by around 50% compared to ChatGPT-4v and by around 40% compared to LLaVA. The promising results demonstrate that KnowLoop can enable the robot to actively ask for human help when the prediction is uncertain.